OpenStudy (anonymous):

how would you integrate the sqrt of (9-x^2) from -3 to 3 . i tried trig sub with x = to 3 sin theta but i get a wrong answer at the end; it get 0

OpenStudy (amistre64):

you will always get a zero.... becuase the left side cancels out the number for the right side

OpenStudy (amistre64):

you either need to move the graph itself to the right by 3 units and adjust the equation; or integrate half of it: from 0 to 3 and then double your result.

OpenStudy (amistre64):

or you could do F(a) - F(b) instead and take the absolute value

OpenStudy (amistre64):

4 - 4 = 0 -4 -(4) = -8 |-8| = 8

OpenStudy (anonymous):

The left side doesnt cancel the right side. It's an even function, you just did the integral wrong.

OpenStudy (anonymous):

no is was suposed to use the power reduction formulas for cos squared

OpenStudy (anonymous):

after you realize that step the problem is solved

OpenStudy (anonymous):

thanks you guys!

OpenStudy (amistre64):

huh... i guess I need to review that part ;) thanx space